A random sample of 16 pigs showed an average increase of weight of 20 kg when a new diet is administered. Assuming

that the increase in weight is normal with variance 4, find 95% confidence interval for average increase in weight due to

new diet. ( ) ( ) )


1
Expert's answer
2021-12-31T04:40:21-0500

The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{x}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(20-1.96\times\dfrac{4}{\sqrt{16}}, 20+1.96\times\dfrac{4}{\sqrt{16}})"

"=(18.04, 21.96)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "18.04 < \mu < 21.96," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(18.04, 21.96)."